OpenStudy (anonymous):

State whether the number is prime or composite. 2 A. prime B. composite @TwilightRain1

OpenStudy (anonymous):

A

OpenStudy (unklerhaukus):

the only factors of a prime number: are itself, and one

OpenStudy (anonymous):

Lot of people viewing it..

OpenStudy (anonymous):

so, can you make 2 by multiplying anything other than itself and 1?

OpenStudy (anonymous):

so how many times can 2 go into 2? 1 and 1 goes into 2...2 times so it is prime

OpenStudy (anonymous):

lol its A because you can only multiply it by one and its self to get i

OpenStudy (anonymous):

I know its A

OpenStudy (anonymous):

good job :P

OpenStudy (anonymous):

State whether the number is prime or composite. 71 A. prime B. composite

OpenStudy (anonymous):

:b

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Wait B

OpenStudy (anonymous):

No A sorry xD

OpenStudy (anonymous):

go with A

OpenStudy (anonymous):

XD its a prime number

OpenStudy (anonymous):

Which example supports the given conjecture? The sum of two even numbers is an even number. A. 1 + 3 = 4 B. 4 + 6 = 10 C. 4 + 5 = 9 D. 0 + 1 = 1

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Which number supports the given conjecture? If a number is divisible by both 2 and 3, then the number is also divisible by 6. A. 16 B. 23 C. 27 D. 36

OpenStudy (anonymous):

D

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

right again

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its the only one that can be divided by 6

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Which equation supports the given conjecture? The cube of an even number is an even number. A. 3 · 2 = 6 B. 23 = 8 C. 53 = 125 D. 4 + 4 + 4 = 12

OpenStudy (unklerhaukus):

use ^'s for powers

OpenStudy (unklerhaukus):

A. 3 · 2 = 6 B. 2^3 = 8 C. 5^3 = 125 D. 4 + 4 + 4 = 12

OpenStudy (unklerhaukus):

^3 is the cube, which options has : {some even number}^3 = {another even number} ?

OpenStudy (anonymous):

(I figured it out